OpenStudy (cloverracer):

construct a rational function that has a vertical asymptote at x = 3 and a removable discontinuity at x = -2.

zepdrix (zepdrix):

Here is a clever way to do it. You want a discontinuity at x=3, so you should have (x-3) showing up somewhere in the function. We need another discontinuity at x=-2, so you should have (x+2) somewhere. When I say "somewhere" I actually mean denominator of the rational function. That's what produces these discontinuities, the denominator.\[\large\rm f(x)=\frac{}{(x-3)(x+2)}\]What should we put in the numerator though? Well, we want this discontinuity at x=-2 to be `removable`. So the factor in the denominator, (x+2), should cancel out to give us this removed value.\[\large\rm f(x)=\frac{(x+2)}{(x-3)(x+2)}\]

zepdrix (zepdrix):

So that's uhh.. ya.. that's one way to do it.

OpenStudy (cloverracer):

okay, makes sense. it doesn't matter if (x+2) is seen twice? due to the fact it's on the numerator and denominator?

zepdrix (zepdrix):

It `must` show up twice. That's what makes it a removable discontinuity. It's sort of a special thing. Usually when you have a removable discontinuity, the factor will be well hidden. Here is an example,\[\large\rm \frac{x^2+3x+2}{x^2+7x+10}\]In this example we have a removable discontinuity at x=-2 but it might not be so obvious at first glance. If we factor the numerator and denominator,\[\large\rm \frac{(x+1)(x+2)}{(x+5)(x+2)}\]you can see both the numerator and denominator have the (x+2), but it was well hidden, ya?

OpenStudy (cloverracer):

ohhh i see now, thank you for the explanation!

zepdrix (zepdrix):

np